Assume that the Section 179 deduction limit is $1,000,000 and that businesses exceeding a total of $2,500,000 in purchases of qualifying personal property will have the Section 179 deduction phase-out dollar-for-dollar. Also assume that the taxpayer has elected out of any bonus depreciation. Please treat each item separately.

1. Equipment in the amount of $300,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $900,000. No other personal or real property was purchased during the year 2. Land in the amount of $30,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $94,000. No other personal or real property was purchased during the year. 3. Equipment in the amount of $10,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $4,000. No other personal or real property was purchased during the year 4. Equipment in the amount of $10,000 was purchased and placed in service during the current year. Net loss before the 179 deduction was $18,000. No other personal or real property was purchased during the year. 5. Equipment in the amount of $1,100,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $1,500,000. No other personal or real property was purchased during the year 6. Equipment in the amount of $2,700,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $940,000. No other personal or real property was purchased during the year 7. Equipment in the amount of $2,700,000 was purchased and placed in service during the current year. Taxable income before the 179 deduction was $750,000. No other personal or real property was purchased during the year 8. An office building in the amount of $1,000,000 was purchased and placed in service during the current year. A fire alarm system in the amount of $100,000 was purchased and installed in the building later in the year. Taxable income before the 179 deduction was $1,500,000 No other personal or real property was purchased during the year. 9. An apartment building in the amount of $500,000 was purchased and placed in service during the current year. A fire alarm system in the amount of $50,000 was purchased and installed in the building later in the year. Taxable income before the 179 deduction was $300,000. No other personal or real property was purchased during the year.
Answer #1

1. taxable income before 179 deduction           = $ 9,00,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 9,00,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

2. taxable income before 179 deduction           = $ 94,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 94,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

3. taxable income before 179 deduction           = $ 4,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 4,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

4. taxable income before 179 deduction           = $ 18,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 18,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

5. taxable income before 179 deduction           = $ 15,00,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 15,00,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

6. taxable income before 179 deduction           = $ 9,40,000

   less: deduction u/s 179                               = $ (10,00,000)

   total income after deduction u/s 179             = $ (60,000)

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is more then $ 25,00,000 so deduction will be there

7. taxable income before 179 deduction           = $ 7,50,000

   less: deduction u/s 179                               = $ (10,00,000)

   total income after deduction u/s 179             = $ (2,50,000)

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

8. taxable income before 179 deduction           = $ 15,00,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 15,00,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there

9. taxable income before 179 deduction           = $ 3,00,000

   less: deduction u/s 179                               = $ 0

   total income after deduction u/s 179             = $ 3,00,000

as the limit for business to purchase personal property should exceed $ 25,00,000 and in this case the limit is less then $ 25,00,000 so no deduction will be there